OpenStudy (anonymous):

Which of the following statements must be true of an equation before you can use the quadratic formula to find the solutions? A. The coefficient of x2 -term can't be 0. B. The coefficient of the x- term must be positive. C. One side of the equation must be 0. D. There can be no term whose degree is higher than 2

OpenStudy (zzr0ck3r):

for \(ax^2+bx+c\) we have \(x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\). So starting with the first option do you see any problem with \(a=0\)

OpenStudy (zzr0ck3r):

?